# Measuring all the noises of LLM Evals

### v1

@@ 3.3 Estimating from samples
Figure 3: Relative errors of the paired variances. $K=5$ samples are drawn from the 500 questions from MATH500, where 1000 real predictions are drawn from models $A$ and $B$ on each question, which is treated as the ground truth. The root mean squared relative errors (rms) of variance components are plotted vs. bootstrap sample size NN, extending beyond the population size 500. The left figure without the correction on $K$ has an unacceptably high 70% relative error even with 2000 data points. The middle figure is with the correction of §3.3 and the right figure uses $1/K$ instead of 1/(K1)1/(K-1).
3.4 Implementation in array notation
In an experiment, we typically evaluate the model on all $N$ questions, and may draw $K$ answers for each question. In this section, we overload our notation with actual samples $A,B\in\mathbb{R}^{N\times K}$ and present the formulas in §3.3 in numpy-style code. For simplicity, we use the same number of predictions $K_{i}=K$ on all questions $x_{i}$. For the unpaired case, we have Table 2. For paired, we have Table 2.
@@ 3.4 Implementation in array notation
3.5 Equivalence to the bootstrap and the sign test
Two other very well-established methods are the bootstrap efron1986bootstrap and paired difference tests such as the sign test dixon1946sign. These methods ask if the observed results are likely when the questions are random for bootstrap, and when the comparison outcomes are random for the sign test. We show they all give the same answer.
The sign test predates bootstrap and instead of randomly resampling the questions, it randomizes the comparison outcomes. It supposes that any difference actually happens with a random probability $\Pr[A(x)\neq B(x)]=0.5$ (null hypothesis) and asks how likely we are to get a more extreme outcome than what we actually observe. This corresponds to a predictor that mixes the predictions of $A$ and $B$ with probability 0.5. If it is quite likely to observe a more extreme outcome under the null hypothesis or equivalently with the mixed predictor, then the conclusion is not statistically significant. Interestingly, the mixed predictor offers a concrete way to get such non-significant differences as long as we have prediction noise.
The key step in both methods is estimating the variance of their respective random procedure. Under the same variance, then the probability to observe the opposite result from an observed mean (bootstrap) is also the same as the probability of observing a more extreme result if the real mean is 0 (sign test). The details are deferred to §B.
4 Experiments
The experimental results of this paper can be found at https://all-the-noises.github.io, where the method is tested on more evals and where the figures are interactive to support more detailed investigations and case studies. We can use this to check that the main conclusion indeed holds on all the evals.
4.1 Exploratory analysis and findings
4.1.1 Data heatmap
To get an overview of our data, we use a heatmap shown in Figure 4. Each row is a different question, sorted from the easiest to the hardest. Each column is a model whose $x$-coordinate is the overall accuracy $\mathrm{E}[A]$ of this model. Like models, we might consider showing questions with $y$-coordinate at their average accuracy instead of rank, but that would depend on which models are included, and less intrinsic to the eval. We immediately observe that models at a similar overall accuracy also does similarly on individual questions. Around 1000 samples are drawn for each question and model for MATH500 in Figure 4. The SWEbench figure is based on the leaderboard data, which had to commit to single answers, but the same overall pattern holds, though with more noise. On MATH500, most of the red regions of bad models are still non-zero.
4.2 All-pairs paired noises and predictable total noise
The noise components can be traded off to some extend so they can depend on the setting and not as predictable as the total. We find at more natural temperatures of $[0.7,1]$, the prediction noise tend to be higher than the data noise across evals. Figure 5 shows the noise components at two different temperatures on CRUXEval, where the breakdown differs but the total noise remains the same.
Figure 5: All noise components on CRUXEval at temperature 0.8 (left) and 0.2 (right). The prediction noise dominates at 0.8 temperature, whereas the data noise dominates at 0.2 temperature, while both still yielding about the same total noise.
Figure 6: Empirical cummulative distribution curves (CDF, solid) and their respective Beta models (dots). Each color represents a different bin by model accuracy. The average accuracy of each question in each bin of models is computed and sorted to produce the CDF curves. The fit is good.
4.2.1 The Beta theory makes good noise predictions and fits the data
Figure 6 shows the the comparison between a $\operatorname{Beta}(a,b)$ distribution and the empirical cumulative distribution (CDF) of the metrics, where the models are binned by range of accuracy. Then $\hat{p}_{1}=c_{1}/K,\ldots,\hat{p}_{N}=c_{N}/K$ is computed for each questions $i$ across the $K$ combined predictions among the binned models, $c_{i}$ is the total number of predictions that evaluated as correct on question $i$. Smaller $a+b$ means more bimodal where $a+b$ tend to decrease when the model accuracy is higher, showing that better models become more bimodal whereas very bad models is usually unimodal. For an extreme example, guessing uniformly on multiple choice questions with $C$ choices means the expected accuracy $u_{i}=1/C$ for all questions, which is has a single mode at $1/C$.
@@ 4.2.1 The Beta theory makes good noise predictions and fits the data
4.3 Exceptions
We note a few exceptions to the general findings. On guessable evals such as MMLU, when the performance is near random chance, the observed noise is closer to the independent prediction rather than the Beta prediction. In Figure 1, swebench-verified, we see 2 outlier results near $x=0.3$ that comes from SWE-Fixer (xie2025swefixer), due to using a deterministic filter instead of drawing another prediction from the model. The predictions of Llama on vLLM is more deterministic than expected and thus has higher data noise than others, likely due to faulty settings. Distilled models tend to have less sharp distribution at the same performance level.
5 Related work
Much work deals with noise in science and machine learning (lehmann2005testing; dror-etal-2018-hitchhikers; hermann2024experimental; card-etal-2020-little, among many others), we adopt the approach of miller2024error to estimate the variance directly. While using samples to estimate the population follows from the law of large numbers, their approach is clarifying, has unique considerations and recommendations for LLMs, and is directly generalizable all metric functions. Their method is also consistent with the bootstrap efron1979bootstrap and the sign test dixon1946sign when applicable (§3.5). efron1986bootstrap clearly described estimating the standard error as the example of not needing bootstrap. The concept of total variance decomposition is present the analysis of variance, but Section 3 of miller2024error applied this concept to LLMs under variance of the conditional mean and mean conditional variance, which we call the prediction noise and the data noise. They also advocated for averaging (resample), but it is not clear how much gains can be expected. In this work, we develop and test the methods of estimating noise types from samples and then empirically measure these variance components on all pairs of models.
For leaderboards, Chatbot arena chiang2024chatbot and CRUXEval gu2024cruxeval computed confidence intervals using bootstrap with a fixed reference model, which allowed them to obtain per-model confidence intervals but is incorrect (Example 4). Furthermore, Chatbot arena aggregates comparisons with all other models to compute their confidence intervals, which necessarily loses any special properties of a particular pair of models. [^1] They no longer provide confidence intervals, but the method is described in their paper. This work actually compares all pairs of models, and use the summary statistics of all the comparisons to get meaningful noise levels.
madaan2024variance measured the noise due to random seeds, which is a form of prediction noise in our view. Any direct approach to measure the noise necessarily ignores the data noise (Example 2). For the data noise, they also proposed using unpaired Bernoulli confidence intervals. This method was adopted by Llama 3 dubey2024llama3 but is much looser than the paired methods (Example 1) and does not separate the data noise and the prediction noise. Improving model ability is hard and expensive while tightening the analysis is well-established.
6 Discussions
6.1 Remarks on the main findings
Predictable total noise.
When the prediction noise is much greater than the data noise, then total noise $\approx$ prediction noise, and prediction noise is predicted by the Beta theory, so the predictable total noise is a consequence of prediction noise $\gg$ data noise. However, Figure 5 shows non-negligible data noise, and where the total noise still follows the prediction even when the data noise is bigger than the prediction noise at a low temperature222except for chain-of-thought predictions, where the prediction noise is still higher. More data is needed to further test if this continue to hold, and if low temperature can reduce prediction noise as the model is training (we guess not, but untested).
prediction noise > data noise.
This means that the model is a bigger source of noise than the data. So methods that tries to model question difficulty such as the Item Response Theory (IRT) or filtering questions polo2024tinybenchmarks should deal with the prediction noise first. Section 5 of madaan2024variance and evalarena find that IRT does not work well. The data filtering method $\tau$- evalarena finds questions that are inversely correlated with the overall model performance. When there is high prediction noise, there is no way to get a reliable signal from a small number of questions or to reliably blame the questions for noise.
6.2 Caveats of the expected metric
@@
With the impressive increase in power, we recommend averaging for most controlled experiments on model architecture, optimization, loss function, etc. that are not expected to have a large effect on the eval and do not target the eval specifically. However, there are some caveats if our goal is to improve a particular eval metric using all the best available methods, and when the independent experiment variable is expected to affect the “sharpness" of the distribution. Example 5 shows how improvement in expected accuracy leads to no increase after aggregation, and how sharpening the distribution on good-enough model can lead to much better expected accuracy. Example 6 show how increased expected accuracy leads to a decrease the final performance when the trade-off is less diversity, which can happen in post-training. Example 7 compares the expected metric vs. alternatives and show that arbitrarily small but consistent changes is significant using the expected metric, clarifying that an independent standard of effect size is also needed.
@@
6.3 Solving hard and special problems
For a problem requiring a long answer where guessing correctly is unlikely, answering even 1 problem might be very significant and interesting. For example, the problem can ask for the proof of an important open problem and the test checks the proof. If a model (or someone) solves such a hard problem then we would not object to the sample size of 1. We probably don’t need to consider this yet. We observe that all pairs of models have enough noisy inconsistencies where model A may beat B on 2 hard questions, but then B beats A on 2 easy questions. If A is so much better than B such that A generated a long and correct solution to an important question, why did it fail on some easy questions where B succeeded? By inspecting the average question level metrics over 1000 samples, we see that worse models also has a smaller chance of guessing the correct answer. For example, if the best model usually answers question $x_{i}$ correctly with $\mathrm{E}_{\epsilon}[A(x_{i})]=0.9$, then we also see $\mathrm{E}_{\epsilon}[B(x_{i})]>0$ detectable by a mere 1000 samples for a much worse model $B$.
In our human experience, we intuitively get a lot of signals by evaluating on hard problems with a number of steps (e.g. interviews, math). The key might be to reflect on the details of how the problem was solved. In contrast, LLM evals and RLVR settings only give back 1 bit back after doing a lot of work. If we continue in the direction of smaller and harder evals, we probably need to also generate more information per question, until then, more questions are needed to satisfy regular statistical standard.
6.4 Applications and Recommendations
In addition to the recommendation of miller2024error, we advocate for more “global" noise measurements to decouple noise analysis from most experiments, so practitioners can focus on their experiments instead of the noise analysis. A motivating example for this recommendation is the Llama 3 report, where many tables show the unpaired standard errors, which is too loose and gives no additional information since they are just a simple function of the accuracy. Since some decoupling is possible, eval and leaderboard builders can naturally conduct the noise analysis by running the all-pairs-paired method to measure all the noise types on their eval, thus taking more responsibility for the statistical reliability of their eval. This can even be done by a third-party as in this work, as long as leaderboard maintainers release their question level results like jimenez2024swebench; jain2024livecodebench, ideally with multiple predictions per question so the noise components can be better estimated. This might be a unique possibility in the LLM paradigm, where both models and evals have global identities to enable this approach.
@@ 6.4 Applications and Recommendations
We have an increasing number of evals of varying quality and size, so we discuss how to make more out of many evals. There are many huge table of results that are hard to interpret in the literature. Averaging MMLU where $N\sim 10k$ and HumanEval where $N\sim 100$ likely leads to worse results than using MMLU alone. Instead of a direct average or the confidence intervals, we advocate using the $z$-score because it is signed and more intuitive, and is well-behaved. Recall the $z$-score $z=\frac{\bar{A}-\bar{B}}{\operatorname{SE}[A-B]}$. This work measured the paired and unpaired versions of prediction, data, and total $\operatorname{SE}$s and is suitable here. Inspecting this list of $z$-scores (one for each eval) and checking their consistency is already informative. To get a meta-$z$-score given the $z$-score of eval $i$, an intuitive method is $\bar{z}=\frac{\sum_{i}w_{i}z_{i}}{(\sum_{i}w_{i}^{2})^{1/2}}$, where some prior beliefs can be incorporated into $w_{i}$. For example, $w_{i}=1$ means that we believe each eval has equal weight, whereas $w_{i}=N_{i}^{1/2}$ means each question has equal weight. While incorporating our prior belief of the importance of each eval into the weight is best suited in LLM evals, $w_{i}=1$ is recommended as a baseline. See zaykin2011weightedz for more related work and analysis, on which we have two remarks specific to LLMs: 1) the effect size of doubling the model size (scaling law) can be a reference on the expected effect size on an eval if desired; 2) giving each question from different evals the same weight is less valid for evals since different evals may capture different abilities, even if this has the most statistical power for a pool of iid questions.
6.5 Limitations
While the method is general, all the empirical data are from correctness evals, so whether the empirical findings generalize to other evals remains to be tested. bowyer2025noclt show that CLT does not work well for less than a few hundred questions, though with 100 questions the errors are only significant in the tails. For simplicity, we did not consider the potentially impactful clustered correction of miller2024error.
Acknowledgments.
I thank Sean O’Brien, Lovish Madaan, Dieuwke Hupkes, Alex Gu, Jiawei Liu, Yuhang Lai, and Sten Sootla for making question-level data available for analysis. I am grateful to Evan Miller, Nicolas Usunier, Zach Rait, Yuxiang Wei, Jannik Kossen, and Ari Holtzman for valuable discussions and feedback; Pedro Rodriguez, Ofir Press, Naman Jain, Baptiste Rozière, Gabriel Synnaeve, Dawn Song, and Zijian Wang for their advice and support. The all-pairs approach is inspired by Chatbot Arena and the clarity of miller2024error greatly helped.
Appendix A Analysis details
A.1 Alternative decomposition
Taking expectation over either the noise or the question first has the same total variance and thus the sums are equal
$\mathrm{Var}_{x,\epsilon}[A]$ $=\mathrm{Var}_{x}[\mathrm{E}_{\epsilon}[A]]+\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A]]$ (25)
$=\mathrm{Var}_{\epsilon}[\mathrm{E}_{x}[A]]+\mathrm{E}_{\epsilon}[\mathrm{Var}_{x}[A]]$ (26)
To see that $\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A(x,\epsilon)]$ indeed corresponds to the prediction noise that we can measure directly,
$\mathrm{Var}_{\epsilon}\left[\frac{1}{N}\sum_{i=1}^{N}A(x_{i},\epsilon_{i})\right]=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}_{\epsilon}[A(x_{i},\epsilon)]\rightarrow\frac{1}{N}\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A(x,\epsilon)].$ (27)
The left hand side is the prediction noise on $N$ particular data points, which converges to the mean if they are iid. One can directly evaluate the prediction noise on a particular dataset by running the eval $K$ times and measure the variance of the $K$ results. That is also correct in expectation but is more noisy since (27) includes equally valid samples not drawn by any particular eval run.
Paired prediction noise from the same model
$=2\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A]]$ (30)
Appendix B Equivalence to bootstrap and sign test
@@
Appendix C Estimator testing
@@
Our main concern is practical accuracy. Tests establish that estimators achieve rms error below 0.25 for $N=100$ sample sizes and below 0.13 for larger samples ($N=400$). We verify this across diverse probability distributions including uniform, Beta, specific distributions and real data. For paired comparisons, tests confirm we get reasonable accuracy (relative rms <0.1<0.1) on the total variance and prediction variance. The data variance is the hardest to estimate since it tend to be the smallest where correcting for the bias in $K$ is critical (Figure 3)..
@@
This testing builds confidence that our estimators behave as predicted and that we understands how to get unbiased estimators in each situation, even as we only recommend two of the simplest estimators.

### v2

@@ A.2 Estimating from samples
Figure 5: Relative errors of the paired variances. $K=5$ samples are drawn from the 500 questions from MATH500, where 1000 real predictions are drawn from models $A$ and $B$ on each question, which is treated as the ground truth. The root mean squared relative errors (rms) of variance components are plotted vs. bootstrap sample size NN, extending beyond the population size 500. The left figure without the correction on $K$ has an unacceptably high 70% relative error even with 2000 data points. The middle figure is with the correction of §A.2 and the right figure uses $1/K$ instead of 1/(K1)1/(K-1).
Same model correction.
$=2\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A]]$ (28)
A.3 Implementation in array notation
In an experiment, we typically evaluate the model on all $N$ questions, and may draw $K$ answers for each question. In this section, we overload our notation with actual samples $A,B\in\mathbb{R}^{N\times K}$ and present the formulas in §A.2 in numpy-style code. For simplicity, we use the same number of predictions $K_{i}=K$ on all questions $x_{i}$. For the unpaired case, we have Table 2. For paired, we have Table 2.
@@ A.3 Implementation in array notation
Appendix B The Beta theory predicts the noise and fits the distribution
Figure 6: Empirical cumulative distribution curves (CDF, solid) and their respective Beta models (dots). Each color represents a different bin by model accuracy. The average accuracy of each question in each bin of models is computed and sorted to produce the CDF curves. The fit is good.
Figure 6 shows the comparison between a $\operatorname{Beta}(a,b)$ distribution and the empirical cumulative distribution (CDF) of the metrics, where the models are binned by range of accuracy. Then $\hat{p}_{1}=c_{1}/K,\ldots,\hat{p}_{N}=c_{N}/K$ is computed for each question $i$ across the $K$ combined predictions among the binned models, $c_{i}$ is the total number of predictions that evaluated as correct on question $i$. Smaller $a+b$ means more bimodal where $a+b$ tend to decrease when the model accuracy is higher, showing that better models become more bimodal whereas very bad models is usually unimodal. For an extreme example, guessing uniformly on multiple choice questions with $C$ choices means the expected accuracy $u_{i}=1/C$ for all questions, which has a single mode at $1/C$.
@@ Appendix B The Beta theory predicts the noise and fits the distribution
Appendix C Caveats and exceptions
C.1 Caveats of the expected metric
@@
With the impressive increase in power, we recommend averaging for most controlled experiments on model architecture, optimization, loss function, etc. that are not expected to have a large effect on the eval and do not target the eval specifically. However, there are some caveats if our goal is to improve a particular eval metric using all the best available methods, and when the independent experiment variable is expected to affect the “sharpness" of the distribution. Example 5 shows how improvement in expected accuracy leads to no increase after aggregation, and how sharpening the distribution on good-enough model can lead to much better expected accuracy. Example 6 shows how increased expected accuracy leads to a decrease the final performance when the trade-off is less diversity, which can happen in post-training. Example 7 compares the expected metric vs. alternatives and shows that arbitrarily small but consistent changes is significant using the expected metric, clarifying that an independent standard of effect size is also needed.
@@
C.2 Exceptions to main findings
We note a few exceptions to the general findings. On guessable evals such as MMLU, when the performance is near random chance, the observed noise is closer to the independent prediction rather than the Beta prediction. In Figure 1, swebench-verified, we see 2 outlier results near $x=0.3$ that come from SWE-Fixer (Xie, 2025), due to using a deterministic filter instead of drawing another prediction from the model. The predictions of Llama on vLLM is more deterministic than expected and thus has higher data noise than others, likely due to faulty settings. Distilled models tend to have less sharp distribution at the same performance level.
Appendix D Applications and Recommendations
In addition to the recommendation of Miller (2024), we advocate for more global noise measurements to decouple noise analysis from most experiments, so practitioners can focus on their experiments instead of the noise analysis. For some motivating examples, in the Llama 3 report many tables show the unpaired standard errors, which is too loose and gives no additional information since they are just a simple function of the accuracy. Merrill et al. (2026) measured the prediction noise (incorrect) and Chan et al. (2025) measured the total noise with different number of seeds (correct but under-specified). We recommend that eval and leaderboard builders run all-pairs-paired method to measure all the noise types on their eval, thus providing well-specified noise measurements and taking responsibility for the statistical reliability of their evals. This can even be done by a third-party as in this work, as long as leaderboard maintainers release their question level results like Jimenez et al. (2024); Jain et al. (2024), ideally with multiple predictions per question so the noise components can be estimated. This might be a unique possibility in the LLM paradigm, where both models and evals have global identities to enable this approach.
@@ Appendix D Applications and Recommendations
We have an increasing number of evals of varying quality and size, so we want to get more signal from more evals. Directly averaging MMLU where $N\sim 10k$ and HumanEval where $N\sim 100$ likely leads to worse results than using MMLU alone, so some weighting is necessary. We advocate using the $z$-score because it is signed and well-behaved. Recall the $z$-score $z=\frac{\bar{A}-\bar{B}}{\operatorname{SE}[A-B]}$. This work measured the paired and unpaired versions of prediction, data, and total $\operatorname{SE}$s and is suitable here. Inspecting this list of $z$-scores, one per eval, and checking their consistency is a good start. To get a meta-$z$-score given the $z$-score of eval $i$, an intuitive method is $\bar{z}=\frac{\sum_{i}w_{i}z_{i}}{(\sum_{i}w_{i}^{2})^{1/2}}$, where some prior beliefs can be incorporated into $w_{i}$. For example, $w_{i}=1$ means that we believe each eval has equal weight, whereas $w_{i}=N_{i}^{1/2}$ means each question has equal weight. While incorporating our prior belief of the importance of each eval into the weight is best suited in LLM evals, $w_{i}=1$ is recommended as a baseline. See Zaykin (2011) for more related work and analysis, on which we have two remarks specific to LLMs: 1) the effect size of increasing the model size (scaling law) can be a reference on the expected effect size on an eval if desired; 2) giving each question from different evals the same weight is less valid for evals since different evals may capture different abilities, even if this has the most statistical power for a pool of iid questions.
Appendix E Equivalence to the bootstrap and the sign test
Two other very well-established methods are the bootstrap Efron and Tibshirani (1986) and paired difference tests such as the sign test Dixon and Mood (1946). These methods ask if the observed results are likely when the questions are random for bootstrap, and when the comparison outcomes are random for the sign test. We show they all give the same answer.
The sign test predates bootstrap and instead of randomly resampling the questions, it randomizes the comparison outcomes. It supposes that any difference actually happens with a random probability $\Pr[A(x)\neq B(x)]=0.5$ (null hypothesis) and asks how likely we are to get a more extreme outcome than what we actually observe. The sign test suggests a mix predictor that mixes the predictions of $A$ and $B$ with probability 0.5, which generates variance from different predictions. Let $\hat{y}$ be the prediction to be evaluated then,
$\hat{y}_{AB}(x)=\begin{cases}\hat{y}_{A}(x)&\text{with prob }1/2\\ \hat{y}_{B}(x)&\text{with prob }1/2.\end{cases}$ (30)
If it is quite likely to observe a more extreme outcome under the null hypothesis or equivalently with the mixed predictor, then the conclusion is not statistically significant.
The key step in both methods is estimating the variance of their respective random procedure. Under the same variance, then the probability to observe the opposite result from an observed mean (bootstrap) is also the same as the probability of observing a more extreme result if the real mean is 0 (sign test). The details follow.
@@ Appendix E Equivalence to the bootstrap and the sign test
Appendix F Estimator testing
@@
Our main concern is practical accuracy. Tests establish that estimators achieve rms error below 0.25 for $N=100$ sample sizes and below 0.13 for larger samples ($N=400$). We verify this across diverse probability distributions including uniform, Beta, specific distributions and real data. For paired comparisons, tests confirm we get reasonable accuracy (relative rms <0.1<0.1) on the total variance and prediction variance. The data variance is the hardest to estimate since it tend to be the smallest where correcting for the bias in $K$ is critical (Figure 5).
@@
This testing builds confidence that our estimators behave as predicted and that we understand how to get unbiased estimators in each situation, even as we only recommend two of the simplest estimators.